The speed of the current in a river is 5mi/hr. It takes 40 min. longer for a girl to paddle 1.2 miles upstream than the distance downstream. What'sher speed in still water?
Answer by Nate(3500)

Rate = x - 5
Time = 40/60 = 2/3
Rate * Time = Distance
(x - 5)(2/3) = 1.2
2x - 10 = 3.6
2x = 13.6
x = 6.8